OpenStudy (saifoo.khan):

To completing square form. Steps are Must. -0.035 x^2+1.4 x+6.

OpenStudy (turingtest):

divide all by -0.035 first...

OpenStudy (saifoo.khan):

x^2 -40x -171.43

OpenStudy (turingtest):

wait, doesn't this have to be equal to zero to do this?

OpenStudy (saifoo.khan):

it is equal to zero.

OpenStudy (turingtest):

-0.035 x^2+1.4 x+6=0 ?????????

OpenStudy (saifoo.khan):

Yes^^

OpenStudy (saifoo.khan):

Hello, Sat.

OpenStudy (anonymous):

no it just has to be the same as what you start with. does not have to be equal to zero

OpenStudy (anonymous):

hello

OpenStudy (turingtest):

but you have to put it in that form at some point, or at least start with knowing what it is equal to.

OpenStudy (anonymous):

if it is equal to zero then to what turing said, divide by the leading coefficient. if it is just an expression you have to factor it out

OpenStudy (saifoo.khan):

it is equal to zero, this is bcz i myself made that equation. So, after divinf i get: x^2 -40x -171.43

OpenStudy (saifoo.khan):

dividing*

OpenStudy (anonymous):

in other words if you start with \[y=-0.035 x^2+1.4 x+6\] you have to end with \[y=\text{ expression} \] that is the same as the initial one

OpenStudy (saifoo.khan):

that is an equation, coz i myself made it.

OpenStudy (saifoo.khan):

BRB

OpenStudy (turingtest):

x^2-40x=171.43 b=-40 completing the square means you add (1/2b)^2 to both sides now so add 0.000156 to both sides at this point I frankly would throw out the idea of completing the square since in this case the decimal thing is getting silly, but theoretically x^2-40x+(1/6400)=(6/0.035)+(1/6400) (x-1/80)^2=(6/0.035)+(1/6400) \[x=1/80+\sqrt{(6/0.035)+(1/6400)}\] I decided to keep fractions here as they are more reasonable than a decimal addition like 0.000156 which would hardly change 171.43 in any noticable way

OpenStudy (anonymous):

\[y=-.035x^2+.14x+6\] \[y=-.035(x^2-4)+6\] \[y=-.035(x^2-4x+4)+6+.035\times 4\] \[y=-.035(x^2-4x+4)+6.14\] \[y=-.035(x-2)^2+6.14\]

OpenStudy (turingtest):

sorry it should be\[1/80 \pm \sqrt{(6/0.035)+(1/6400)}\] if y=0

OpenStudy (anonymous):

the method i used above is just rewriting it as you would any quadratic you wanted to change from the form \[y=ax^2+bx+c\] into \[y=a(x+h)^2+k\] i certainly would not use it to find the zeros, although you could

OpenStudy (anonymous):

in this form you can see that the vertex is \[(2,6.14)\] but you could find the vertex by computing \[-\frac{b}{2a}\]as well

OpenStudy (anonymous):

in fact that is a shorter way of completing the square. find \[-\frac{b}{2a}\] in your case it is 2. the write as \[y=-.035(x-2)^2+\text{ something }\] and you find the "something" by replacing x by 2 in the original expression and see what you get

OpenStudy (turingtest):

Ah, I see satellite's method is more general and efficient than mine, but for the problem posted with y=0 am I not correct in my method too?

OpenStudy (anonymous):

sure if it is set equal to zero you do not have to keep it the same and are free to divide by the leading coefficient

OpenStudy (anonymous):

if \[ax^2+bx+c=0\iff x^2+\frac{b}{a}x+\frac{c}{a}=0\]

OpenStudy (anonymous):

but if you need to end with what you start with, then you have to "factor it out" and cannot divide by it. because \[ax^2+bx+c=0\neq x^2+\frac{b}{a}x+\frac{c}{a}=0\]

OpenStudy (saifoo.khan):

that is incorrect form which sat gave me! :/